Please answer all

A. Carbohydrates contain 4 kcalories per gram, protein contains 4 kcalories per gram, and fat contains 9 kcalories per gram. To calculate the percentage of kcalories from protein, carbohydrate or fat, (1) determine the total kcalories in the food (calculate the kcals of protein, carbohydrate and fat and add them together) and (2) to determine the percentages of each divide the kcalories in the individual component (protein, carbohydrate or fat) by the total kcalories in the food, and multiply by 100.

Calculate the energy (total kcalories) in a bean and cheese burrito which contains 55 gm carbohydrate, 15 gm protein and 12 gm fat. Calculate the percentage of protein, carbohydrate and fat in the burrito. Show your work.

B. Carbohydrates contain 4 kcalories per gram, protein contains 4 kcalories per gram, and fat contains 9 kcalories per gram. To calculate the percentage of kcalories from protein, carbohydrate or fat, (1) determine the total kcalories in the food (calculate the kcals of protein, carbohydrate and fat and add them together) and (2) to determine the percentages of each divide the kcalories in the individual component (protein, carbohydrate or fat) by the total kcalories in the food and multiply by 100.

Calculate the energy from a meal containing 44 gm protein, 115 gm carbohydrate and 30 gm fat. Calculate the percentage of protein, carbohydrate and fat in the meal. Show your work.

C. Carbohydrates contain 4 kcalories per gram, protein contains 4 kcalories per gram, and fat contains 9 kcalories per gram. To calculate the grams of protein, carbohydrate or fat, (1) multiply the total kcals of the food by the percentage of each (protein, carb, fat) to get the percentage of kcals (protein, carb, fat). Then divide by the kcals/gram to get the grams.

Calculate the grams of protein, carbohydrate and fat in a meal that is 850 calories and is 25% protein, 45% carbohydrate and 30% fat. Show your work.
Answer #1

A) Calculate total calories

Calories from carbohydrates - 55gm

Each calorie contains 4 grams of carbohydrates.

Total calories from carbohydrate = 55/4 = 13.75 kcal

calories from protein - 15gm

Each calorie contain 4 grams of protein

Total calories from protein = 15 / 4 = 3.75 kcal

Calories from fat - 12gm.

Each calorie contains 9 grams of fat.

Total calories from fat = 12 / 9 = 1.33 kcal

Total calories in bean and cheese burrito = 13.75 + 3.75+ 1.33 = 18.83.

Calculating the total percentage

Percentage of kcal = Percentage of individual component/total kcal x 100

Percentage of kcal from carbohydrates = 13.75 / 18.83 x 100

= 73.02%

Percentage of kcal from protein = 3.75 / 18.83 x 100

= 19.91%

Percentage of Kcal from fat = 1.33 / 18.83 x 100

= 7.06%

B) Calculate the total calories of the meal:

Carbohydrate - 115gm

Calories from carbohydrate = 115 / 4 = 28.75kcal.

protein = 44gm

Calories from protein = 44 /4 = 11kcal

fat - 30gm

Calories from fat = 30 / 9 = 3.33 kcal

Total kcal in a meal = 28.75 + 11 +3.33

Total kcal in a meal = 43.08 kcal

Percentage of kcal from carbohydrates = 28.75 / 43.08 x 100

Percentage of kcal from carbohydrates = 66.73%

Percentage of kcal from protein = 11 / 43.08 x 100

Percentage of kcal from protein = 25.53%   

Percentage of kcal from fat = 3.33 / 43.08 x 100

Percentage of kcal from fat = 7.72%

C) Calculate the grams of protein, carbohydrates, and fat

Total calories - 850 calories

Protein percentage - 25% or 0.25

Calculating grams of individual component = total calories x percent / calories per gram

Calculating grams of protein = 850 x 0.25 / 4

Grams of protein = 53.12gm

Carbohydrate percentage = 45% or 0.45

Calculating grams of carbohydrates = 850 x 0.45 / 4

Grams of carbohydrate = 95.62gm

fat percentage = 30%or 0.30

Calculating grams of fat = 850 x 0.30 / 9

Grams of fat = 28.33gm